OpenStudy (anonymous):

Solve the inequality and graph its solution: 5n – 10 > 25

OpenStudy (whpalmer4):

First, find the solution if you had an equality instead of an inequality. \[5n -10 = 25\]\[n =\] Next, examine the inequality sign. Does it contain an =? (\(\le,\ge\)) If so, you use a closed circle or filled dot, because the endpoint is included. Otherwise, use an open dot or empty circle, because the endpoint is not included. Finally, decide which portion to shade. I test with 0 or (0,0) because the arithmetic is easy. If we set \(n=0\), is the original inequality true? If so, shade the portion that includes \(n=0\); otherwise, shade the other part.

OpenStudy (anonymous):

thank you